OpenStudy (anonymous):

A man stands on the roof of a 19.0 m-tall building and throws a rock with a velocity of magnitude 30.0 m/s at an angle of 34.0 above the horizontal. You can ignore air resistance. Calculate the magnitude of the velocity of the rock just before it strikes the ground.

OpenStudy (anonymous):

There are a couple ways to do this; do you prefer kinematics or conservation of energy methods?

OpenStudy (anonymous):

They both reduce to the same equation in the end: \[(v_f)^2-(v_o)^2 = 2a(y_f-y_o)\]

OpenStudy (anonymous):

kinematics

OpenStudy (anonymous):

I still need help

OpenStudy (anonymous):

Ok. Do you recognize the formula I put there?

OpenStudy (anonymous):

It can be derived either by combining the first two fundamental kinematics equations, \[x_f=x_o+v_o(t)+\frac{1}{2}a(t)^2 \space \space and \space \space v_f=a(t)\] or by using conservation of energy, i.e. kinetic + potential before = kinetic + potential after. Either way you can solve for final velocity with \[(v_f)^2−(v_o)^2=2a(y_f−y_o) \rightarrow v_f=\sqrt{(v_o)^2+2a(y_f−y_o)}\]

OpenStudy (anonymous):

And since it only wants the magnitude of the velocity, you don't have to worry about the vectors (except in the positive vs. negative sense).

OpenStudy (anonymous):

did you get 36.2 as the answer?

OpenStudy (anonymous):

(sorry for the delay, was working on a different problem) (I haven't worked it out yet, so let me go grab my calculator..)

OpenStudy (anonymous):

I got 35.68, so you probably just have some round-off error in there somewhere.